Question:

If ๐‘Ž and ๐‘ are two sides of a right angled hypotenuse ๐‘ and let ๐‘ be the perpendicular from the opposite

vertex on the hypotenuse, show that:

1. (๐›ฟ ๐‘/๐›ฟ๐‘Ž)b =b3/c3

2. (๐›ฟ๐‘/๐›ฟ๐‘Ž )A=b/c




1
Expert's answer
2021-08-23T09:06:19-0400

Answer:-

1. From Pythagoras theorem



"c^2=a^2+b^2\\"


The area of the triangle



"A=\dfrac{1}{2}ab=\dfrac{1}{2}pc"

Then



"p=\dfrac{ab}{c}""p=p(a, b)=\dfrac{ab}{\sqrt{a^2+b^2}}"

Differentiate byย "a"



"\dfrac{\partial p}{\partial a}=\dfrac{b}{\sqrt{a^2+b^2}}-\dfrac{a^2b}{\sqrt{a^2+b^2}(a^2+b^2)}""=\dfrac{a^2b+b^3-a^2b}{\sqrt{a^2+b^2}(a^2+b^2)}"

Substituteย "c=\sqrt{a^2+b^2}"



"\dfrac{\partial p}{\partial a}=\dfrac{b^3}{c^3}"

2.



"A=const=>\dfrac{1}{2}ab=const"

Letย "a=kb, k>0."ย Thenย "c^2=a^2+b^2=k^2b^2+b^2=b^2(1+k^2)."



"\dfrac{b}{c}=\sqrt{\dfrac{1}{k^2}}=const""p=\dfrac{b}{c}(a)"

Use thatย "\dfrac{b}{c}=const,"ย ifย "A=const."ย Then



"(\dfrac{\partial p}{\partial a})_A=\dfrac{b}{c}"